OpenStudy (anonymous):

How many three digits numbers are there such that sum of the digits is even, number is odd and divisible by 5? @Directrix @JamesJ

OpenStudy (anonymous):

Just checking, my answer is 45.

OpenStudy (anonymous):

Hey Ishaan, try it took me 2 mints but my friends think I am wrong :(

OpenStudy (anonymous):

Actually she says that there were no option like 45 in the actual test, I have a feeling that she didn't remember the question properly.

OpenStudy (anonymous):

sum of the digits is even, numbers are odd and the sum is divisible by 5 or sum of the digits is even, numbers are odd and numbers are divisible by 5

OpenStudy (anonymous):

according to her version, "sum of the digits is even, numbers are odd and numbers are divisible by 5"

OpenStudy (anonymous):

Could you give me one example? I am getting confused :-/

OpenStudy (anonymous):

btw, for "sum of the digits is even, numbers are odd and the sum is divisible by 5" is still 45

OpenStudy (anonymous):

Okay, 215 235 255 275 295 305 325 345 365 385 415 435 455 475 495 505 525 545 565 585

OpenStudy (anonymous):

Okay I wrote a brute-force and seems like I am right! :)

OpenStudy (anonymous):

For the case you suggested the numbers are: 109 127 145 163 181 217 235 253 271 299 307 325 343 361 389 415 433 451 479 497 505 523 541 569 587 613 631 659 677 695 703 721 749 767 785 811 839 857 875 893 901 929 947 965 983

OpenStudy (anonymous):

Lol Thanks, I thought individual digits must be odd, which made the question impossible. I overthink things :-/

OpenStudy (anonymous):

Over-thinking is good, and don't worry I am a Fool myself lol

OpenStudy (anonymous):

I am getting 90 in total (Inclusion of numbers ending with ..0 and ..5). So, I think 45 for Odd.

OpenStudy (ash2326):

@FoolForMath Last digit should be 5 to be divisible by 5 If first digit is even, second digit should be odd= \(4\times 5\times 1=20\) If first digit is odd, second digit should be even=\(5\times 5\times \times 1=25\) Total three no. digits which satisfy the conditions= 45

OpenStudy (phi):

How many three digits numbers are there such that sum of the digits is even, number is odd and divisible by 5? 45 assuming no leading zeros in the 3 digit number. in each decade, there is only 1 number divisible by 5 i.e. ends in 5 in each hundred, there are 10 decades, 5 of which have digits that sum to an even # so 5 numbers per hundred. there are 9 hundreds (100, 200,...900), so 9*5= 45

OpenStudy (jamesj):

I get 45 as well.

OpenStudy (asnaseer):

the last digit can be 0 or 5 for the number to be divisible by 5. e.g. 110 130 150 ... so I get 90?

OpenStudy (asnaseer):

oh - the number must be odd - so it must end in 5 - therefore only 45